OpenStudy (anonymous):

3x + 2y = 5 5x + 2y = 7 Based on the given system of equations, which of the following is not true? i got 2x = 2??

OpenStudy (saifoo.khan):

what does it mean by "which of the following is not true?

OpenStudy (anonymous):

its multiple choice question

OpenStudy (saifoo.khan):

so what u have to find?

OpenStudy (anonymous):

here.... 3x + 2y = 5 5x + 2y = 7 Based on the given system of equations, which of the following is not true? A.2x = 2 B.8x = 12 C.4y = 4

OpenStudy (saifoo.khan):

B.8x = 12

OpenStudy (saifoo.khan):

this bcz i got the coordinates (1, 1)

OpenStudy (anonymous):

The coordinates are (1, 1) 2x = 2 => x = 1 => 3.1 + 2y = 5 => 2y = 2 => y = 1